OpenStudy (anonymous):

Solve y over negative 2 + 5 = 13. -16 16 -36 36

OpenStudy (nincompoop):

y/(-2+5) = 13?

OpenStudy (shamil98):

no nin

OpenStudy (shamil98):

\[\huge \frac{ y }{ -2 } + 5 =13\]

OpenStudy (nincompoop):

y/-(2+5) = 13?

OpenStudy (shamil98):

pretty sure its the format i posted it in

OpenStudy (nincompoop):

I am not sure until it is clarified. I am not in the business of certainty.

OpenStudy (anonymous):

the format shamil posted it in is my question

OpenStudy (shamil98):

use the business of logic :)

OpenStudy (shamil98):

anyways subtract 5 from both sides what do you get?

OpenStudy (anonymous):

y-2 and 8 ?

OpenStudy (anonymous):

oppisite of division so then i multiply it by -2x8 ?

OpenStudy (shamil98):

yeah

OpenStudy (shamil98):

y = -2*8 y = ?